OpenStudy (anonymous):

factor the algebraic expression below in terms of a single trigonometric function. cos x - sin 2^x - 1

OpenStudy (ghazi):

sin^2 x-1= cos^2 x... \[-cosx- (\cos^2 x-1)-1= -\cos^2 x-\cos x=0\]

OpenStudy (ghazi):

now solve for x

OpenStudy (ghazi):

just consider cos x= y ...\[y^{2}-y=0...y(y-1)=0\]

OpenStudy (anonymous):

thank you !!!!!

OpenStudy (ghazi):

:))